OpenStudy (anonymous):

Find the sum of the series and explain how you know the sum converges to this value. Sum of (((-1)^(n+1))/n) n=1, infinty.

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} ((-1)^{n+1} x^n)/n \]

OpenStudy (blockcolder):

That's ln(1+x).

OpenStudy (anonymous):

ln(1+1)=ln(2) but how you know it converges to this value?

OpenStudy (blockcolder):

Alternating series test.

OpenStudy (anonymous):

Are you familiar with Taylor Series or Maclaurin Series?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If you try to compute the Maclaurin series of:\[\ln(x+1)\]you will obtain the power series in your question.

OpenStudy (zarkon):

differentiate the series...sun the then geometric series...integrate the series

OpenStudy (zarkon):

*integrate the closed form after summing the geometric